solve differential equation thru laplace transfrom and inverse laplace

dx + 2 d.²x - dx = 2X= 4 + et dt at² dt

with complete solution
Answer #1

Am) Given differential Equation is dex dt3 +2.138 dx • 26 2x ute df2 df Take Lapalace Transferm 83x(s) + 25% x(s) - S x(s) –Go) AE syls 1. SSD A = (sxo) - 8 은 -2 2 (0) (0) () (아 B- (S-D (s) S-2. 0 - S2 -8 10-8 -) 12. Q (2) [21)(2) 2X3M c. (9) 79 11E = (S+2) (5) \ S=-2 81- 낸 S아) -8 (-)(--) (-2) (24) 8x3 TD - 3 So, A - -2, 33 능 C=) 2 D 13 - , 1 0 9 E- E = -3 So, X(s) 13 +Solution is provided in uploaded pics.

If I helped you then Support me with a Thumps up